OpenStudy (annie96):

Question 1.1. Solve. 5y = 135 (Points : 1) –19 –27 27 19 Question 2.2. Solve. -1/8c=2 (Points : 1) 16 4 –4 –16

OpenStudy (shadowlegendx):

So is this solving for the variable?

OpenStudy (shadowlegendx):

Cuz thats what I'm guessing

OpenStudy (annie96):

yes

OpenStudy (shadowlegendx):

So the problem 5y = 135 is basically \[5 \times y = 135\]

OpenStudy (shadowlegendx):

So 5 multiplied by what equals 135?

OpenStudy (annie96):

27

OpenStudy (shadowlegendx):

Correct

OpenStudy (shadowlegendx):

So the answer for question #1 Is 27

OpenStudy (shadowlegendx):

Now for #2

OpenStudy (yttrium):

c = -1/8

OpenStudy (shadowlegendx):

The problem -1/8c=2 is basically \[\frac{ 1 }{ 8 } \times c = 2\]

OpenStudy (shadowlegendx):

Sorry thats supposed to be a negative 1/8

OpenStudy (shadowlegendx):

So what is c?

OpenStudy (shadowlegendx):

This is a similar process to the other problem

OpenStudy (annie96):

the answer is 4

OpenStudy (shadowlegendx):

Correct :)

OpenStudy (annie96):

Thank you so much

OpenStudy (shadowlegendx):

\(\bbox[5pt, #33fd33 ,border:5px solid #872af0]{\bf\huge\ Good \ Job!!!}\)

OpenStudy (shadowlegendx):

Thanks Annie :)